[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jun 26, 2017

Craig Mattocks - NOAA Federal craig.mattocks at noaa.gov
Mon Jun 26 15:45:23 UTC 2017


Update on TAFB item #1:

NCO Windows Sysadmin Marc Laurent determined that the noise from the Situational Awareness PC was caused by a failing graphics card. A new card has been ordered and will be shipped overnight for installation tomorrow.

Craig

> On Jun 26, 2017, at 11:31 AM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:
> 
> Here is today's summary of NHC computer operations: 
> --- Matt Sardi
> 
> ----------
> 
> TAFB:
> 1. Situational Awareness PC was making loud noise behind the Atlc desk. Sounds like the fan. Turned it off overnight and the monitors as well.
> 
> 2. Upper monitor of SHINER was blinking due to loose cable; Helpdesk ticket opened, NCO resolved.
> 
> HSU:
> 1. Dave Z's verification page is not working for the total 2017 eastern pacific season-- storm by storm it seems to be, but the cumulative graphic is just showing the first storm. Dave will fix today (need to remove old "BAM" models).
> 
> 2. SHIPS-RI is still broken due to the lack of satellite input. According to Mark DeMaria, this is because the GOES-W NH sector we get does not go far enough east. When a TC gets west of about 103 deg W, there is enough data in the current sector for the RII to run. 
> 
> So far this year, nearly all (maybe all) of the EP TCs have been too far east for the RII to run. Mark is working on a fix to use the GOES-east sector instead of GOES-west for EP TCs east of 105 W, so the RII should run for all the EP TCs. Unfortunately, that won't be implemented until after he gets back from AL.
> 
> 3. SAB satellite fix failed to import; only e-mail and NESDIS web page versions were available. TSB will investigate if this remains a problem.
> 
> Other items:
> 1. No GOES-16 imagery available since 20Z Sunday in AWIPS/N-AWIPS. TSB is investigating.
> 
> 2. NCO will be performing a bandwidth upgrade on Wednesday, 6/26; network service degradation/interruptions are possible during two 20 minute windows (1530-1550Z & 1620-1640Z). TAFB/HSU service backup may be required if outages last longer than expected.
> 
> NHC RFCs:
> 1. RFC 239 - Update the N-AWIPS workstation FVWM (F virtual window manager) config files for the TAFB operational accounts (surface, atlmar, pacmar) to pipe legacy Gempak script output to the /usr/bin/tee command and write to log files. The legacy TAFB Gempak scripts do no generate log files, which makes it challenging to track down problems. This change will facilitate troubleshooting. Was to be implemented on June 19 at 1500Z; postponed due to CWD; new implementation date is Monday, June 26 at 1500Z (11 am EDT).
> 
> 2. RFC 241 - Implement several enhancements to the TAFB 'low bandwidth' graphical marine forecast webpages, including addition of intermediate 12 hourly time steps, image labeling and web viewer improvements, domain adjustments, etc. Modify sfcprog launch and surface analysis web to create new 12 hourly/intermediate low bandwidth web images on adjusted forecast sub-domains; update javascript viewer on NHC website to view forecast grid images. Was to be implemented on June 20 at 1600Z; postponed due to CWD; new implementation date is Monday, June 26 at 1600Z (noon EDT).
> 
> 3. RFC 243 - When TCs are upgraded from "Potential Tropical Cyclone" to "Tropical Storm" an additional check in the code is required so that the existing wind radii information can be used. This check is added in this RFC.
> 
> 




More information about the Ncep.nhc.nco_contacts mailing list